OpenStudy (anonymous):

use the following functions: f(x) = x + 5 g(x) = x2 -4x + 5 h(x) = 3x2 - 2x + 5 find: f [g(2)]

OpenStudy (lin.ivory):

g(2) = (2)^2 - 4(2) +5 = 1 f(g(2))= f(1) = 6 therefore f(g(2)) = 6

OpenStudy (anonymous):

wheres the 2 come from for the x values in the g equation?

OpenStudy (anonymous):

g(2) will be the input of f(x) so x of f(x) will be g(2)